OpenStudy (anonymous):

Factor completely: 9x^2 + 42x + 49

OpenStudy (anonymous):

use the quadratic formula \[-b \pm \sqrt{b^2 - 4ac}\div2a\] where a = 9 b=42 c=49.. Just plug in the numbers and simplify.

OpenStudy (riley):

9x^2 +42x+49 So let's break it down. (3x+7) (3x+7) = 9x^2+42x+49 So your answer will be: (3x+7)^2

OpenStudy (riley):

Yep what Tyler said, using the quadratic formula often times helps A LOT when your working with more complicated equations. But when it's more simpler stuff like this, you can just factor out the equation. You just have to break it down and find like values that plug in and work out.

OpenStudy (anonymous):

The Quadratic Equation will work for every quadratic problem where as factoring like Riley did will only work for the answers that are non-imaginary. Both ways are right but there are some problems where you will have to use the Quadratic Formula